# Activated LoRA: Fine-tuned LLMs for Intrinsics

### v4

@@ 2.1 Training
Our implementation of aLoRA is in the supplementary material. It seamlessly supports standard Huggingface training libraries Wolf et al. (2019) as well as inference/generation methods (with or without using available base model KV cache) if needed (e.g. for testing or proofs of concept when the efficiencies of vLLM (Zhu et al., 2023) are not needed). Later in this section, we will also show inference experiments using our modification of the more efficient (SOTA) inference package vLLM to support aLoRA.
@@ 3 Inference timing results on vLLM
Figure 3: Comparison of aLoRA and LoRA when used as evaluators in a simple agentic pattern. Top left: Multiplicative speedup of an aLoRA evaluator vs LoRA, showing up to 35x improvement depending on base model and prompt length. Top right: Multiplicative speedup for the end-to-end pipeline including the base model generation (256 tokens) and 1 or 5 parallel eval (adapter) generations (16 tokens each). Despite the large fixed cost of the base model call, end-to-end aLoRA speedups are still significant, highlighting LoRA inefficiency. Bottom row: Log-log plots for the wall clock evaluation, showing that even for small models, the delay for LoRA becomes meaningful in absolute terms as the prompt and number of evaluations in the agentic pipeline scale.
@@
To prove this experimentally, we modified vLLM (Zhu et al., 2023) to be able to perform inference on aLoRAs with base model prefix cache reuse, and compared LoRA and aLoRA inference on 7 small ($\leq 14$B) LLMs in the setting where the initial base model prompt length is varied, the base model generated answer is comprised of 256 tokens, and each adapter evaluation is given by 16 tokens. Results are in Figure 3, showing timing for both 1 evaluator and 5 evaluators. [^5] While 5 evaluators may initially seem a lot, this is a stand-in for more complex pipelines where the input, retrieved documents, and generation are all being checked, as in Danilevsky et al. (2025).
The speedup factor of aLoRA increases as prompt length increases, adapter generation time improving over 2 to 7$\times$ even for 250-token prompts, and over 20$\times$ for most models on the longest prompts. Despite most of the compute advantages coming from the prefill savings (time to 1st token of the adapter), we still see highly meaningful speedups in the end-to-end pipeline which also includes the large fixed cost of generating 256 tokens with the base model. Finally, in the last row, despite the adapters only generating 16 tokens, we see that the LoRA delays are noticeable in wall clock time. Note these wall clock delays will be further magnified for more complex agentic or reasoning pipelines where many rounds of input and generation happen before the overall pipeline completes.
@@ 4 Finetuning Accuracy Results
Having demonstrated significant inference time speedups for aLoRA, it remains to ascertain whether we lose generation accuracy or quality. We first train LoRA and aLoRA adapters on 4 LLMs on a set of benchmark SFT tasks, and then consider a set of more challenging “intrinsics”-type tasks for which well-engineered, very recent LoRA adapters already exist against which we can compare.
@@ 4.1 Benchmark SFT tasks
In Brüel-Gabrielsson et al. (2024), a collection of 1000 instruction SFT tasks were curated and LoRA adapters were trained for each. These tasks were drawn from from the Super-Natural Instructions Wang et al. (2022a) benchmark collection of datasets, which in turn drew from sources such as MMLU Hendrycks et al. (2021) etc.
From this list, we selected 7 tasks at random, roughly split between multiple-choice and freeform outputs. We excluded from consideration any tasks with (a) too small datasets, (b) overly open-ended instructions, or (c) difficulty such that trained adapters did no better than random chance. The selected tasks and datasets are detailed in Appendix F. For each task, we trained both LoRA and aLoRA adapters for 4 instruct-tuned models of various sizes. To ensure a fair comparison, for each task-model pair, using the validation set performance we performed grid search over 4 rank values and 5 learning rate values [^6] Note that this experiment represents 1120 separate adapter training runs. and followed typical LoRA SFT best practices for all parameters. The invocation sequence was set to equal the model’s chat template generation prompt. See Appendix F for further details.
@@ 4.2 Intrinsics Tasks
In this subsection, we consider real-world tasks that fit the framework of intrinsics, specifically, settings where the task requires a potentially long input that either (a) already has base model KV cache available from either being input to or generated by the base model, or (b) does not have (complete) base model KV cache available, but base model KV cache created through an aLoRA call will be used in a subsequent base model call. We draw from those proposed in Danilevsky et al. (2025) for a RAG pipeline, though these intrinsics are not limited to RAG. See Danilevsky et al. (2025) for deeper motivation for each intrinsic and extensive experimental results comparing their LoRA adapters to strong baselines.
In Danilevsky et al. (2025), LoRA adapters were trained for the intrinsics using the Granite 3.2 8b Instruct model. For direct comparison to their released adapters, we train corresponding aLoRA adapters (using the same datasets and chat-template-based formatting) and compare their performance in terms of generation quality and/or accuracy. [^7] As they released LoRA adapters for only one base model, we limit ourselves to their choice of base model. Note that since all the below models use the same base model, they can be swapped in and out as needed in the same flow. The reader can envision the possibilities of what can be built with these intrinsics, e.g. for RAG. See the App. G for additional details for this section.
@@
Uncertainty Quantification: This intrinsic provides a Certainty score for model responses to user questions. The model will respond with a number from 0 to 9, corresponding to $5\%,15\%,25\%,\dotsc,95\%$ confidence respectively. Training data for these confidence scores are obtained by applying the UQ calibration method of Shen et al. (2024a) to a large, diverse set of benchmark question answering datasets, quantizing the predicted confidences and using these predicted confidences as SFT targets for the adapter. Following the chat template, the invocation sequence is 4 tokens: <|start_of_role|>certainty<|end_of_role|>. Note that the model is evaluating responses from its base model - in other words, it cannot be applied to generations from other models. The aLoRA architecture is thus particularly well-suited to this use case. In practice, the goal is to provide a highly efficient uncertainty score without having to resort to expensive larger judge models.
@@
Answerability Determination: This intrinsic assesses whether a user’s final query in a multi-turn conversation can be answered given the retrieved documents. In RAG or other settings, this decides whether to proceed with generation or abstain with an “I don’t know” response. The invocation sequence is set to <|start_of_role|>answerability<|end_of_role|>. We tested on binary answerability classification on the single-turn SQUADRun Benchmark Rajpurkar et al. (2018) with the user query and the supporting documents, and the multi-turn MT-RAG Benchmark Katsis et al. (2025) using full multi-turn conversation history along with the supporting documents. Figure 6 shows the results. Overall, the aLoRA model does not lose performance relative to the LoRA model.
@@
Retrieval recall evaluation (Recall@k) with different query rewrite strategies, evaluated on full, non-standalone and standalone subsets of MT-RAG dataset Katsis et al. (2025) are shown in Figure 1(a). All retrieved passages are obtained using the Elser retriever with the same settings as in the above paper. We evaluate on three different subsets of MT-RAG detailed in the appendix.
Answer quality evaluation using RAGAS Faithfulness (RAGAS-F) and RAD-Bench on full, non-standalone and standalone subsets of MT-RAG dataset are shown in Table 1(b) (see appendix for details). Note that throughout, performance numbers for aLoRA and LoRA are within a point or two. See Danilevsky et al. (2025) for additional comparisons showing that these approaches outperform benchmarks and are very close to the performance with gold rewrites.
@@
Jailbreak Detection: This intrinsic is designed for detecting jailbreak risk within user prompts. Prompts with jailbreak risk vary across a wide range of attack styles - from direct instructions, to encoding-style, social-hacking based attacks and even ones that exploit special token or context overload (Rawat et al., 2024). In our experiments we focused on training intrinsics for detecting social hacking style of adversarial prompts. As with prior intrinsics, the aLoRA/LoRA detectors are trained in the same conditions with a rank of 32. The intrinsic is trained to return a binary label - “Y” indicating jailbreak risk present and “N” indicating no risk.
After training, we evaluate the jailbreak intrinsic on new out-of-distribution datasets not used in training to robustly assess generalizability Zizzo et al. (2025). The out-of-distribution datasets comprise of 3,282 samples containing a mixture of samples with jailbreak risk (Shen et al., 2024b; Lin et al., 2023; Wan et al., 2024) and benign samples (Conover et al., 2023). As with the other intrinsics, we see very little performance difference between LoRA and aLoRA.
@@ 5 Conclusion
We presented Activated LoRA (aLoRA), a novel modification of the LoRA framework enabling efficient and dynamic adaptation of large language models without requiring recomputation of the key-value (KV) cache. By modifying LoRA to adapt weights only for tokens generated after activation, aLoRA facilitates seamless integration into multiturn settings, enabling the use of specialized “intrinsics” for well-defined operations within a broader conversational or processing pipeline.
@@ Appendix C Proofs
For the causal decoder-only transformers we consider, the keys and values (actually, all internal states) prior to $t_{\mathrm{invoke}}$ are identical for the base model and any aLoRA adapter model using (6). Specifically, $K^{base}_{1:t_{\mathrm{invoke}}-1}=K^{adapter}_{1:t_{\mathrm{invoke}}-1}$, and $V^{base}_{1:t_{\mathrm{invoke}}-1}=V^{adapter}_{1:t_{\mathrm{invoke}}-1}$. Inference with the aLoRA adapted model can be done causally one token at a time (by simply increasing $t$ in (6) iteratively) with KV cache reuse.
@@
Since the transformer’s subsequent layer outputs (including all MLP and layer-norm states) are deterministic functions of these $Q,K,V$ up to $i$, it follows by induction on layer depth that all internal states for tokens $11,\dots,t_{\mathrm{invoke}}-$ 1 agree between the two models. In particular the cached key- and value-matrices satisfy
@@
During generation at time $t\geq t_{\mathrm{invoke}}$, both models proceed token-by-token via exactly the same causal-attention mechanism (1), merely appending the new $(Q_{t},K_{t},V_{t})$ row and reusing the previously cached rows. Since up to $t-1$ those rows coincide, the adapter may reuse the base model’s KV cache for the first $t_{\mathrm{invoke}}-1$ tokens, and then continue to grow its own cache for tokens $t_{\mathrm{invoke}},\dots,t-1$. This establishes that aLoRA inference can indeed be performed causally, one token at a time, with full KV cache reuse. ∎
@@
Consider invoking an adapter with $T_{cache}$ tokens of input for which a base model KV cache exists and $T_{new}\ll T_{cache}$ input tokens without cache. [^8] We presume that $T_{new}$ is the invocation sequence and following. The first token generated by the aLoRA adapter requires $O(T_{cache}T_{new})$ operations, while LoRA requires $O((T_{cache}+T_{new})^{2})$. Furthermore, aLoRA must maintain only $O(T_{new})$ additional KV cache memory, while LoRA requires additional $O(T_{cache}+T_{new})$. If $N$ distinct adapters share the first $T_{cache}$ tokens as input, aLoRA costs become $O(NT_{cache}T_{new})$ and $O(NT_{new})$, while LoRA has $O(N(T_{cache}+T_{new})^{2})$ and $O(N(T_{cache}+T_{new}))$.
@@
Let the input be partitioned into $T_{\mathrm{cache}}$ tokens whose base-model cache is already available, and $T_{\mathrm{new}}$ tokens on which the adapter must act without precomputed cache. We compare the cost of generating the first new token:
@@
      form the attention scores by multiplying these new queries against the cached-plus-new key-matrix of size $(T_{\mathrm{cache}}+T_{\mathrm{new}})\times d_{k}$, costing $O((T_{\mathrm{cache}}+T_{\mathrm{new}})\,d_{k})$,
@@
If $N$ distinct adapters each process the same $T_{\mathrm{cache}}+T_{\mathrm{new}}$ context but disjoint $T_{\mathrm{new}}$ segments, the aLoRA per-adapter cost scales as $O(T_{\mathrm{cache}}+N\,T_{\mathrm{new}}T_{\mathrm{cache}})$ time and $O(N\,T_{\mathrm{new}})$ memory, whereas LoRA’s cost scales on the full quadratic context for each adapter (totaling $O\bigl{(}N\,(T_{\mathrm{cache}}+T_{\mathrm{new}})^{2}\bigr{)}$ time and $O\bigl{(}N\,(T_{\mathrm{cache}}+T_{\mathrm{new}})\bigr{)}$ memory). ∎
@@ F.1 Models
Llama 3.2 1B Instruct (meta-llama/Llama-3.2-1B-Instruct on Huggingface), Llama 3.2 3B Instruct (meta-llama/Llama-3.2-3B-Instruct), Llama3.1 8B (meta-llama/Llama-3.1-8B-Instruct) Grattafiori et al. [2024], and Mistral 7B (mistralai/Mistral-7B-Instruct-v0.3) Singh et al. [2023].

### v5

@@ 2.1 Training
Our implementation of aLoRA is in the supplementary material. It seamlessly supports standard Huggingface training libraries [43] as well as inference/generation methods (with or without using available base model KV cache) if needed (e.g. for testing or proofs of concept when the efficiencies of vLLM [45] are not needed). Later in this section, we will also show inference experiments using our modification of the more efficient (SOTA) inference package vLLM to support aLoRA.
@@ 3 Inference timing results on vLLM
Figure 3: Comparison of aLoRA and LoRA when used as evaluators in a simple agentic pattern. Top left: Multiplicative speedup of an aLoRA evaluator vs LoRA, showing up to 35x improvement depending on base model and prompt length. Top right: Multiplicative speedup for the end-to-end pipeline including the base model generation (256 tokens) and 1 or 5 parallel eval (adapter) generations (16 tokens each). Despite the large fixed cost of the base model call, end-to-end aLoRA speedups are still significant, highlighting LoRA inefficiency. Bottom row: Log-log plots for the wall clock evaluation, showing that even for small models, the delay for LoRA becomes significant in absolute terms as the prompt and number of evaluations in the agentic pipeline scale.
@@
To prove this experimentally, we modified vLLM [45] to be able to perform inference on aLoRAs with base model prefix cache reuse, and compared LoRA and aLoRA inference on 7 small ($\leq 14$B) LLMs in the setting where the initial base model prompt length is varied, the base model generated answer is comprised of 256 tokens, and each adapter evaluation is given by 16 tokens. [^7] Code for our vLLM implementation can be found at https://github.com/tdoublep/vllm/tree/alora. Results are in Figure 3, showing timing for both 1 evaluator and 5 evaluators. [^8] While 5 evaluators may initially seem a lot, this is a stand-in for more complex pipelines where the input, retrieved documents, and generation are all being checked, as in [5]. In this experiment, we set the LoRA to have rank 8 and the aLoRA to be rank 32, illustrating that any need for rank increase with aLoRA has negligible inference time effect since the adapter parameters are still typically much less than 1% of the base model parameters.
[…]
@@ 4 Finetuning Accuracy Results
Having demonstrated significant inference time speedups for aLoRA, it remains to ascertain whether we lose generation accuracy or quality. We first train LoRA and aLoRA adapters on 4 LLMs on a set of benchmark SFT tasks, and then consider a set of more challenging “intrinsics”-type tasks for which well-engineered, very recent LoRA adapters already exist against which we can compare. Our code implementing Activated LoRA has been contributed to the Huggingface PEFT library https://github.com/huggingface/peft [25], which we also use for LoRA. [^9] PEFT contains the model implementation and handles inference. Training in our experiments is done using a standard Huggingface TRL [38] trainer (SFTTrainer).
@@ 4.1 Benchmark SFT tasks
In [2], a collection of 1000 instruction SFT tasks were curated and LoRA adapters were trained for each. These tasks were drawn from the Super-Natural Instructions [42] benchmark collection of datasets, which in turn drew from sources such as MMLU [11] etc.
From this list, we selected 7 tasks at random, roughly split between multiple-choice and freeform outputs. We excluded from consideration any tasks with (a) too small datasets, (b) overly open-ended instructions, or (c) difficulty such that trained adapters did no better than random chance. The selected tasks and datasets are detailed in Appendix F. For each task, we trained both LoRA and aLoRA adapters for 4 instruct-tuned models of various sizes. To ensure a fair comparison, for each task-model pair, using the validation set performance we performed grid search over 4 rank values and 5 learning rate values [^10] Note that this experiment represents 1120 separate adapter training runs., and followed typical LoRA SFT best practices for all parameters. Note that we did not observe any pattern to the performance-maximizing rank value in these experiments, in particular there was not a clear pattern of LoRA or aLoRA needing higher or lower rank. The invocation sequence was set to equal the model’s chat template generation prompt. See Appendix F for further details.
@@ 4.2 Intrinsics Tasks
In this subsection, we consider real-world tasks that fit the framework of intrinsics, specifically, settings where the task requires a potentially long input that either (a) already has base model KV cache available from either being input to or generated by the base model, or (b) does not have (complete) base model KV cache available, but base model KV cache created through an aLoRA call will be used in a subsequent base model call. We draw from those proposed in [5] for a RAG pipeline, though these intrinsics are not limited to RAG. See [5] for deeper motivation for each intrinsic and extensive experimental results comparing their LoRA adapters to strong baselines.
In [5], LoRA adapters were trained for the intrinsics using the Granite 3.2 8b Instruct model. For direct comparison to their released adapters, we train corresponding aLoRA adapters (using the same datasets and chat-template-based formatting) and compare their performance in terms of generation quality and/or accuracy. [^11] As they released LoRA adapters for only one base model, we limit ourselves to their choice of base model. Note that since all the below models use the same base model, they can be swapped in and out as needed in the same flow. The reader can envision the wide range of possibilities enabled by these intrinsics, e.g. for RAG. See the Appendix G for additional details for this section.
Uncertainty Quantification: This intrinsic provides a Certainty score for model responses to user questions. The model will respond with a number from 0 to 9, corresponding to $5\%,15\%,25\%,\dotsc,95\%$ confidence respectively. Training data for these confidence scores are obtained by applying the UQ calibration method of [32] to a large, diverse set of benchmark question answering datasets, quantizing the predicted confidences and using these predicted confidences as SFT targets for the adapter. Following the chat template, the invocation sequence is 4 tokens: <|start_of_role|>certainty<|end_of_role|>. Note that the model is evaluating responses from its base model - in other words, it cannot be applied to generations from other models. The aLoRA architecture is thus particularly well-suited to this use case. In practice, the goal is to provide a highly efficient uncertainty score without having to resort to expensive larger judge models.
@@
Answerability Determination: This intrinsic assesses whether a user’s final query in a multi-turn conversation can be answered given the retrieved documents. In RAG or other settings, this decides whether to proceed with generation or abstain with an “I don’t know” response. The invocation sequence is set to <|start_of_role|>answerability<|end_of_role|>. We tested on binary answerability classification on the single-turn SQUADRun Benchmark [30] with the user query and the supporting documents, and the multi-turn MT-RAG Benchmark [15] using full multi-turn conversation history along with the supporting documents. Figure 6 shows the results. Overall, the aLoRA model does not lose performance relative to the LoRA model.
@@
Retrieval recall evaluation (Recall@k) with different query rewrite strategies, evaluated on full, non-standalone and standalone subsets of MT-RAG dataset [15] are shown in Figure 1(a). All retrieved passages are obtained using the Elser retriever with the same settings as in the above paper. We evaluate on three different subsets of MT-RAG detailed in the appendix.
Answer quality evaluation using RAGAS Faithfulness (RAGAS-F) and RAD-Bench on full, non-standalone and standalone subsets of MT-RAG dataset are shown in Table 1(b) (see appendix for details). Note that throughout, performance numbers for aLoRA and LoRA are within a point or two. See [5] for additional comparisons showing that these approaches outperform benchmarks and are very close to the performance with gold rewrites.
@@
Jailbreak Detection: This intrinsic is designed for detecting jailbreak risk within user prompts. Prompts with jailbreak risk vary across a wide range of attack styles - from direct instructions, to encoding-style, social-hacking based attacks and even ones that exploit special token or context overload [31]. In our experiments we focused on training intrinsics for detecting social hacking style of adversarial prompts. As with prior intrinsics, the aLoRA/LoRA detectors are trained in the same conditions with a rank of 32. The intrinsic is trained to return a binary label - “Y” indicating jailbreak risk present and “N” indicating no risk.
After training, we evaluate the jailbreak intrinsic on new out-of-distribution datasets not used in training to robustly assess generalizability [46]. The out-of-distribution datasets comprise of 3,282 samples containing a mixture of samples with jailbreak risk [33, 20, 39] and benign samples [4]. As with the other intrinsics, we see very little performance difference between LoRA and aLoRA.
@@ 5 Conclusion
We presented Activated LoRA (aLoRA), a novel modification of the LoRA framework enabling efficient and dynamic adaptation of large language models without requiring recomputation of the key-value (KV) cache. By modifying LoRA to adapt weights only for tokens generated after activation, aLoRA facilitates seamless integration into multiturn settings, enabling the use of specialized “intrinsics” for well-defined operations within a broader conversational or processing pipeline. While aLoRA adapters sometimes require higher rank $r$ than LoRA, this is not a meaningful downside at inference time since the rank $r$ matrix multiplications are a very small part of overall inference costs (although training costs do increase).
@@ Appendix C Proofs
For the causal decoder-only transformers we consider, the keys and values (actually, all internal states) prior to $t_{\mathrm{invoke}}$ are identical for the base model and any aLoRA adapter model using (6). Specifically, $K^{\mathrm{base}}_{1:t_{\mathrm{invoke}}-1}=K^{\mathrm{adapter}}_{1:t_{\mathrm{invoke}}-1}$, and $V^{\mathrm{base}}_{1:t_{\mathrm{invoke}}-1}=V^{\mathrm{adapter}}_{1:t_{\mathrm{invoke}}-1}$. Inference with the aLoRA adapted model can be done causally one token at a time (by simply increasing $t$ in (6) iteratively) with KV cache reuse.
@@
Since the transformer’s subsequent layer outputs (including all MLP and layer‐norm states) are deterministic functions of these $Q,K,V$ up to $i$, it follows by induction on layer depth that all internal states for tokens $1,\dots,t_{\mathrm{invoke}}-1$ agree between the two models. In particular the cached key‐ and value‐matrices satisfy
@@
During generation at time $t\geq t_{\mathrm{invoke}}$, both models proceed token‐by‐token via exactly the same causal‐attention mechanism (1), merely appending the new $(Q_{t},K_{t},V_{t})$ row and reusing the previously cached rows. Since up to t1t-1 those rows coincide, the adapter may reuse the base model’s KV cache for the first $t_{\mathrm{invoke}}-1$ tokens, and then continue to grow its own cache for tokens $t_{\mathrm{invoke}},\dots,t-1$. This establishes that aLoRA inference can indeed be performed causally, one token at a time, with full KV cache reuse. ∎
@@
Consider invoking an adapter with $T_{\mathrm{cache}}$ tokens of input for which a base model KV cache exists and $T_{\mathrm{new}}\ll T_{\mathrm{cache}}$ input tokens without cache. [^12] We presume that $T_{\mathrm{new}}$ is the invocation sequence and following. The first token generated by the aLoRA adapter requires $O(T_{\mathrm{cache}}T_{\mathrm{new}})$ operations, while LoRA requires $O((T_{\mathrm{cache}}+T_{\mathrm{new}})^{2})$. Furthermore, aLoRA must maintain only $O(T_{\mathrm{new}})$ additional KV cache memory, while LoRA requires additional $O(T_{\mathrm{cache}}+T_{\mathrm{new}})$. If $N$ distinct adapters share the first $T_{\mathrm{cache}}$ tokens as input, aLoRA costs become $O(NT_{\mathrm{cache}}T_{\mathrm{new}})$ and $O(NT_{\mathrm{new}})$, while LoRA has $O(N(T_{\mathrm{cache}}+T_{\mathrm{new}})^{2})$ and $O(N(T_{\mathrm{cache}}+T_{\mathrm{new}}))$.
@@
Let the input be partitioned into $T_{\mathrm{cache}}$ tokens whose base‐model cache is already available, and $T_{\mathrm{new}}$ tokens on which the adapter must act without precomputed cache. We compare the cost of generating the first new token:
@@
      form the attention scores by multiplying these new queries against the cached-plus-new key‐matrix of size $(T_{\mathrm{cache}}+T_{\mathrm{new}})\times d_{k}$, costing $O((T_{\mathrm{cache}}+T_{\mathrm{new}})\,d_{k})$,
@@
If $N$ distinct adapters each process the same $T_{\mathrm{cache}}+T_{\mathrm{new}}$ context but disjoint $T_{\mathrm{new}}$ segments, the aLoRA per‐adapter cost scales as $O(T_{\mathrm{cache}}+N\,T_{\mathrm{new}}T_{\mathrm{cache}})$ time and $O(N\,T_{\mathrm{new}})$ memory, whereas LoRA’s cost scales on the full quadratic context for each adapter (totaling $O\bigl(N\,(T_{\mathrm{cache}}+T_{\mathrm{new}})^{2}\bigr)$ time and $O\bigl(N\,(T_{\mathrm{cache}}+T_{\mathrm{new}})\bigr)$ memory). ∎
@@ F.1 Models
Llama 3.2 1B Instruct (meta-llama/Llama-3.2-1B-Instruct on Huggingface), Llama 3.2 3B Instruct (meta-llama/Llama-3.2-3B-Instruct), Llama3.1 8B (meta-llama/Llama-3.1-8B-Instruct) [9], and Mistral 7B (mistralai/Mistral-7B-Instruct-v0.3) [35].